OpenStudy (anonymous):

solve for x 6/x +14 = 4/3x

OpenStudy (anonymous):

6/x + 14 = 4/3x Let 1/x=y so eq becomes 6y+14=4y/3 6y-4y/3=-14 14y/3=-14 y=-3 put 1/x=-3 so x=-1/3

OpenStudy (anonymous):

I got \[(21 + \sqrt{513}) / 4 \approx 10.9123\]

OpenStudy (anonymous):

i dont think this (21 + sqrt(513))/4 is the solution of this problem

OpenStudy (anonymous):

well I was looking at the problem as 6/x + 14 = (4/3)x

OpenStudy (anonymous):

6/x+14=4/3x 18+42x=4x 42x-4x=-18 38x=-18 x=-18/38 x=-9/18

OpenStudy (anonymous):

no -9/18 is also not the solution

OpenStudy (anonymous):

right approach is by taking 1/x=y still if u want to solve it the way agradecompetent has done then 6/x+14=4/3x 2nd step is wrong, it will be 18x+42x^2=4x and when u solve it further u'll get x=-1/3

OpenStudy (anonymous):

@ agradecompetent : hope u agree with me